# FEDONet : Fourier-Embedded DeepONet for Spectrally Accurate Operator Learning

### v1

@@ 3.6 Allen–Cahn Equation
We evaluate the performance of the proposed Fourier-Embedded DeepONet (FEDONet) in modeling the one-dimensional Allen–Cahn equation—a nonlinear reaction-diffusion system widely used to describe phase separation processes in multi-component systems. The governing partial differential equation is given by:
@@
Figure 15 shows qualitative predictions for the best-performing test case. FEDONet sharply resolves the interface zones and phase transitions, while DeepONet produces smoother approximations with notable blurring near discontinuities. The spectral inductive bias introduced by randomized Fourier features enables FEDONet to retain high-frequency details crucial to phase-field dynamics.
Figure 16: Energy Spectrum and Pointwise absolute error across space and time for a representative sample.
The pointwise absolute error plot in Figure 16 emphasizes the localized improvements of FEDONet in capturing steep interfaces. DeepONet accumulates substantial errors in regions with high spatial gradients, whereas FEDONet maintains predictive accuracy throughout the domain, suggesting better handling of dissipative nonlinearities.
Figure 16 compares the spatial energy spectra of predicted and true solutions. FEDONet accurately tracks the ground truth spectrum across all wavenumbers, preserving the high-frequency tail associated with sharp phase boundaries. In contrast, DeepONet suffers from spectral attenuation at $k>20$, revealing an inability to fully capture dissipative and oscillatory dynamics inherent in the Allen–Cahn system.
These results collectively demonstrate that the incorporation of spectral priors via Fourier embeddings in FEDONet enhances both pointwise accuracy and spectral fidelity. The improved resolution of phase interfaces and better retention of high-frequency information position FEDONet as a robust surrogate model for nonlinear, multiscale PDEs such as the Allen–Cahn equation.
@@ 3.7 Kuramoto–Sivashinsky Equation
Figure 17: Reconstruction of a median-performing KS trajectory. Relative $L^{2}$ error: FEDONet = 16.36%, DeepONet = 89.63%.
Figure 17 shows the reconstructed spatiotemporal field for a median test case. The FEDONet model effectively captures the nonlinear evolution of the solution, preserving the wave envelope, phase dynamics, and intermittent modulation typical of KS attractors. In stark contrast, the vanilla DeepONet produces incoherent and smoothed reconstructions, with pronounced phase lag and energy collapse beyond $t>20$. This degradation is reflected in a high relative error of $89.63\%$ compared to just $16.36\%$ for FEDONet.
In Figure 18, we highlight the best-performing test case. FEDONet achieves a relative $L^{2}$ error of only $10.77\%$, accurately modeling both the fine-scale chaotic filaments and large-scale rolling structures. Notably, the trajectory remains consistent with the ground truth over long integration times, demonstrating that the spectral trunk can stabilize predictions and suppress drift. Such performance underscores the model’s ability to retain energy in high-frequency modes—essential for modeling non-periodic fine-grained behavior that typifies chaotic fields.
@@
Figure 19: Worst-performing test sample. Even in failure cases, FEDONet remains qualitatively reasonable. Relative $L^{2}$ error: FEDONet = $58.02\%$, versus near-complete divergence for DeepONet.
@@
Figure 20: Training convergence comparison. FEDONet achieves faster and more stable optimization due to implicit spectral regularization.
Figure 21: Distribution of relative $L^{2}$ errors across 512 unseen test samples. FEDONet maintains lower error in both mean and variance, demonstrating improved robustness and generalization in chaotic regimes.
@@
In totality, the Kuramoto–Sivashinsky benchmark clearly illustrates the limitations of standard DeepONet trunks in handling chaotic, broadband dynamics. The introduction of Fourier embeddings into the trunk network provides an effective inductive bias that promotes spectral alignment, temporal coherence, and long-horizon accuracy. FEDONet emerges as a compelling model for operator learning in complex regimes, particularly those characterized by sensitive dependence on initial conditions and multiscale feature interactions.
@@ 4 Summary
In this work, we propose the Fourier-Embedded DeepONet (FEDONet) as a principled extension to the classical DeepONet architecture for learning nonlinear operators governed by partial differential equations (PDEs). By injecting randomized Fourier features into the trunk input, FEDONet expands the hypothesis space of DeepONets, offering enhanced spectral representation, improved numerical conditioning, and faster convergence during training—all without increasing runtime complexity. FEDONet consistently outperforms the vanilla DeepONet. Notably, it delivers dramatic reductions in relative $\ell_{2}$ error—most prominently in multiscale and high-frequency scenarios such as the Kuramoto–Sivashinsky equation (76% $\rightarrow$ 21%).
@@
Table 2: Qualitative performance across seven benchmark problems.
Method 2-D Poisson 1-D Burgers Lorenz–63 Lorenz–96 Allen–Cahn KS Eikonal
DeepONet
FEDONet ✓✓ ✓✓ ✓✓ ✓✓
  • Legend: ✓ decent;  ✓✓ excellent;  ✗ sub-par.
Beyond pointwise accuracy, FEDONet preserves physical consistency across modalities. Energy spectrum analyses reveal that FEDONet better captures inertial and dissipation range behavior, avoiding the artificial spectral pile-up observed in DeepONets. In geometry-sensitive tasks like the Eikonal equation, it reconstructs sharp features and discontinuities with higher fidelity.
Taken together, these results establish FEDONet as a robust, scalable, and architecture-agnostic enhancement for neural operator learning. The proposed spectral trunk serves as a drop-in replacement that aligns the inductive biases of DeepONets with the spectral demands of complex PDE systems, providing a promising path forward for efficient surrogate modeling in scientific machine learning.
@@ 5 Future Work
Another avenue involves making the Fourier embedding parameters learnable. While the current design uses fixed random frequency matrices, enabling adaptive tuning of these frequencies could allow the model to better align with the intrinsic spectral content of the data. This modification could yield a more expressive and data-efficient hypothesis space while preserving the training stability benefits of approximate whitening. Finally, incorporating uncertainty quantification and physics-informed constraints into the FEDONet framework—via Bayesian formulations or PINO-style residual losses—may further enhance robustness, interpretability, and generalization in scientific computing tasks involving sparse or noisy data.

### v2

@@ 3.6 Allen–Cahn Equation
We evaluate the performance of the proposed Fourier-Embedded DeepONet (FEDONet) in modeling the one-dimensional Allen–Cahn equation nonlinear reaction-diffusion system widely used to describe phase separation processes in multi-component systems. The governing partial differential equation is given by:
@@
Figure 15 shows qualitative predictions for the best-performing test case. FEDONet sharply resolves the interface zones and phase transitions, while DeepONet produces smoother approximations with notable blurring near discontinuities. The spectral inductive bias introduced by randomized Fourier features enables FEDONet to retain high-frequency details crucial to phase-field dynamics. The pointwise absolute error plot in Figure 16 emphasizes the localized improvements of FEDONet in capturing steep interfaces. DeepONet accumulates substantial errors in regions with high spatial gradients, whereas FEDONet maintains predictive accuracy throughout the domain, suggesting better handling of dissipative nonlinearities.
Figure 16: Allen-Cahn - Energy Spectrum and Pointwise absolute error across space and time for a representative sample.
Figure 16 compares the spatial energy spectra of predicted and true solutions. FEDONet accurately tracks the ground truth spectrum across all wavenumbers, preserving the high-frequency tail associated with sharp phase boundaries. In contrast, DeepONet suffers from spectral attenuation at $k>20$, revealing an inability to fully capture dissipative and oscillatory dynamics inherent in the Allen–Cahn system. These results collectively demonstrate that the incorporation of spectral priors via Fourier embeddings in FEDONet enhances both pointwise accuracy and spectral fidelity. The improved resolution of phase interfaces and better retention of high-frequency information position FEDONet as a robust surrogate model for nonlinear, multiscale PDEs such as the Allen–Cahn equation.
@@ 3.7 Kuramoto–Sivashinsky Equation
Figure 17: Median-performing KS trajectory. Relative $L^{2}$ error: FEDONet = 16.36%, DeepONet = 89.63%.
Figure 17 shows the reconstructed spatiotemporal field for a median test case. The FEDONet model effectively captures the nonlinear evolution of the solution, preserving the wave envelope, phase dynamics, and intermittent modulation typical of KS attractors. In stark contrast, the vanilla DeepONet produces incoherent and smoothed reconstructions, with pronounced phase lag and energy collapse. This degradation is reflected in a high relative error of $89.63\%$ compared to just $16.36\%$ for FEDONet.
In Figure 18, we highlight the best-performing test case. FEDONet achieves a relative $L^{2}$ error of only $10.77\%$, accurately modeling both the fine-scale chaotic filaments and large-scale rolling structures. Notably, the trajectory remains consistent with the ground truth over long integration times, demonstrating that the spectral trunk can stabilize predictions and suppress drift. Such performance underscores the model’s ability to retain energy in high-frequency modes essential for modeling non-periodic fine-grained behavior that typifies chaotic fields.
@@
Figure 19: Worst-performing KS test sample. Even in failure cases, FEDONet remains qualitatively reasonable. Relative $L^{2}$ error: FEDONet = $58.02\%$, versus near-complete divergence for DeepONet.
@@
Figure 20: KS Training convergence comparison. FEDONet achieves faster and more stable optimization due to implicit spectral regularization.
Figure 21: Distribution of relative $L^{2}$ errors across 512 unseen test samples for KS Equation. FEDONet maintains lower error in both mean and variance, demonstrating improved robustness and generalization in chaotic regimes.
@@
In totality, the Kuramoto-Sivashinsky benchmark clearly illustrates the limitations of standard DeepONet trunks in handling chaotic, broadband dynamics. The introduction of Fourier embeddings into the trunk network provides an effective inductive bias that promotes spectral alignment, temporal coherence, and long-horizon accuracy. FEDONet emerges as a compelling model for operator learning in complex regimes, particularly those characterized by sensitive dependence on initial conditions and multiscale feature interactions.
@@ 4 Summary
In this work, we propose the Fourier-Embedded DeepONet (FEDONet) as a principled extension to the classical DeepONet architecture for learning nonlinear operators governed by partial differential equations (PDEs). By injecting randomized Fourier features into the trunk input, FEDONet expands the hypothesis space of DeepONets, offering enhanced spectral representation, improved numerical conditioning, and faster convergence during training all without increasing runtime complexity. FEDONet consistently outperforms the vanilla DeepONet. Notably, it delivers dramatic reductions in relative $\ell_{2}$ error, most prominently in multiscale and high-frequency scenarios such as the Kuramoto-Sivashinsky equation (76% $\rightarrow$ 21%).
@@
Beyond pointwise accuracy, FEDONet preserves physical consistency across modalities. Energy spectrum analyses reveal that FEDONet better captures inertial and dissipation range behavior, avoiding the artificial spectral pile-up observed in DeepONets. In geometry-sensitive tasks like the Eikonal equation, it reconstructs sharp features and discontinuities with higher fidelity. Taken together, these results establish FEDONet as a robust, scalable, and architecture-agnostic enhancement for neural operator learning. The proposed spectral trunk serves as a drop-in replacement that aligns the inductive biases of DeepONets with the spectral demands of complex PDE systems, providing a promising path forward for efficient surrogate modeling in scientific machine learning.
@@ 5 Future Work
Another avenue involves making the Fourier embedding parameters learnable. While the current design uses fixed random frequency matrices, enabling adaptive tuning of these frequencies could allow the model to better align with the intrinsic spectral content of the data. This modification could yield a more expressive and data-efficient hypothesis space while preserving the training stability benefits of approximate whitening. Finally, incorporating uncertainty quantification and physics-informed constraints into the FEDONet framework via Bayesian formulations or PINO style residual losses may further enhance robustness, interpretability, and generalization in scientific computing tasks involving sparse or noisy data.